### v1

@@ 5.2 Semi-implicit GePUP-ES-SDIRK schemes
In light of Theorems 9 and 10, one way to preserve the monotonic decrease of the kinetic energy and the exponential decay of the divergence residue is to employ an algebraically stable RK method in GePUP-ES-RK. Gauss-Legendre RK methods are algebraically stable and have a minimal number of stages for a given temporal order of accuracy. However, their employment in GePUP-ES-RK necessitates either the coupling of all intermediate stage values of $\mathbf{w}$ or the use of complex arithmetic. Thus we turn to singly diagonal implicit RK (SDIRK) methods that satisfy
@@
Definition 7.
A GePUP-ES-SDIRK scheme is a GePUP-ES-RK scheme in which an algebraically stable SDIRK method is employed as the RK method and in which stage values for each intermediate stage $i=1,\ldots,s$ are decomposed as
@@
    where $\widetilde{\mathbf{w}}^{(i)}$ is an approximation of $\mathbf{w}^{(i)}$ obtained by (57) and (EAU-1,2).
  3. 2.
@@
    $\left\{\begin{array}[]{rl}(1-\nu\gamma k\Delta)\mathbf{w}_{0}^{(i)}\par&=\mathbf{w}^{n}+k\sum\limits_{j=1}^{i}a_{i,j}\left(\mathbf{g}^{(j)}-\nabla q_{1}^{(j)}\right)+\nu k\sum\limits_{j=1}^{i-1}a_{i,j}\Delta\mathbf{w}_{0}^{(j)};\\ (1-\nu\gamma k\Delta)\mathbf{w}_{\ell}^{(i)}&=-ka_{i,\ell}\left(\widetilde{\mathbf{u}}^{(\ell)}\cdot\nabla\widetilde{\mathbf{u}}^{(\ell)}+\nabla q_{2}^{(\ell)}\right)+\nu k\sum\limits_{j=\ell}^{i-1}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)}.\end{array}\right.$
  5. 4.
@@
The above steps (GES.1–6) are direct consequences of the GePUP-ES-RK scheme (56)–(55), the property (67) of SDIRK, and the decomposition (68).
Substitute the third decomposition in (68) into (56a,b), separate the terms with and without $r^{(i)}$, notice that $r^{(i)}$ is a scalar, and we have (GES.1–2). In step (GES.1), we decouple $q^{(i)}$ from $\mathbf{w}^{(i)}$ by approximating $\mathbf{w}^{(i)}$ with $\widetilde{\mathbf{w}}^{(i)}$ using the same method for calculating $\widetilde{\mathbf{u}}^{(i)}$.
For SDIRK, the upper bound of the summation in (56d) is $i$. Substitute (68) into (56c,d), separate the terms with and without $r^{(i)}$, switch the summation order
$\begin{array}[]{l}\sum_{j=1}^{i}a_{i,j}\sum_{\ell=1}^{j}r^{(\ell)}\Delta\mathbf{w}_{\ell}^{(j)}=\sum_{\ell=1}^{i}r^{(\ell)}\sum_{j=\ell}^{i}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)},\end{array}$
and we have (GES.3). Then (GES.4) follows from (56g).
(GES.5) follows from substituting the second decomposition in (68) into (56f),
@@
and moving the $r^{(i)}$-term on the RHS to the LHS. Finally, the parenthesis on the LHS of (GES.5) must be no less than 1 because of (53), Lemma 9, and the inner product of $\mathbf{u}_{i}^{(i)}$ to the equation on $\mathbf{w}_{i}^{(i)}$ in (GES.3).
(55) and (GES.1–6) are the complete algorithmic steps of GePUP-ES-SDIRK. Theorems 9 and 10 state that both the velocity divergence and the modified energy in GePUP-ES-SDIRK decrease monotonically provided that the employed SDIRK is algebraically stable; an example is given in (69).
In this work, we discretize the continuous spatial operators in Definition 7 by the fourth-order collocated finite-volume operators in [5, Sec. 3 & 4] to obtain a fully discrete GePUP-ES-SDIRK scheme. However, it is emphasized that what we have proposed is not a single scheme but a space of solvers, each of which can be easily constructed by making menu choices for ‘orthogonal’ policies that span the solver space. The orthogonal structure of this solver space is very conducive to reusing in a black-box manner the legacy of classical finite volume/difference methods and the wealth of theory and algorithms for numerically solving ODEs; see Table 1. It is the GePUP-E formulation that makes this black-box reuse possible.
It would be ideal if the conclusions of Theorems 9 and 10 could also hold in the fully discrete case. Unfortunately, for the finite volume discretization on collocated grids, the fully discrete counterparts of Theorems 9 and 10 only hold asymptotically, i.e., in the limit of $k$ and $h$ simultaneously approaching zero. We defer to future research the investigation of suitable spatial discretizations so that Theorems 9 and 10 also hold in the fully discrete case.
@@ 6 Tests
In this section, we test a particular fully discrete GePUP-ES-SDIRK scheme by several numerical experiments to confirm the analytic results in previous sections. We employ the fourth-order finite-volume discretizations in [5] and a fourth-order, algebraically stable SDIRK method proposed by Du, Ju & Lu [36],
$\gamma$&$\gamma$&&\\ $\frac{1}{2}$&$\frac{1}{2}-\gamma$&$\gamma$&\\ $1-\gamma$&$2\gamma$&$1-4\gamma$&$\gamma$\\ \hline\cr&$\mu$&$1-2\mu$&$\mu$\end{tabular},\qquad\left\{\begin{array}[]{l}\gamma=\frac{1}{\sqrt{3}}\cos\frac{\pi}{18}+\frac{1}{2},\\ \mu=\frac{1}{6(2\gamma-1)^{2}}.\end{array}\right.$ (69)
Table 1: Main orthogonal policies of fully discrete GePUP-ES-RKs.
Orthogonal policies Our choice Other possible options
@@
variable location collocated (as in [5]) staggered
@@
in Definitions 6 and 7 (57) & (EAU-1,2) both in time and in space
Along with other possible options of the major orthogonal policies, we show in Table 1 our choices that constitute the particular GePUP-ES-SDIRK scheme to be tested. According to the analysis in Section 5, this particular INSE solver should be high-order accurate both in time and in space, preserve monotonic decrease of the total kinetic energy, and have an asymptotically vanishing velocity divergence.
In all tests, numerical implementations are performed on a rectangular domain in Cartesian coordinates and thus the electric boundary conditions in (39b) are enforced by setting homogeneous Dirichlet conditions for the tangential velocity components and a homogeneous Neumann condition for the normal velocity. The initial condition of cell-averaged velocity is calculated by Boole’s rule, a sixth-order formula of Newton-Cotes quadrature. Since exact solutions are unavailable, we define the computational error via Richardson extrapolation, i.e., by the difference of the solution on the current grid and that of the next finer grid.
For different values of the penalty parameter $\lambda=$1, 10, and 100, error norms resulting in all our numerical tests remain the same up to the first two significant digits; this differs from the results reported in [22, 23]. This is not surprising since projecting $\mathbf{w}$ to $\mathbf{u}$ with $\mathbf{u}\cdot\mathbf{n}=0$ in (39c,d) and setting $\mathbf{w}^{0}=\mathbf{u}^{0}$ in (40) already imply that $|\mathbf{n}\cdot\mathbf{w}|$ be small. Nonetheless, a positive-valued $\lambda$ guarantees the no-penetration condition be fulfilled. Hereafter in this section we only show results in the case of $\lambda=1$.
@@ 6.1 Single-vortex tests
(a) $t$=40
(b) $t$=60
Figure 2: Snapshots of vorticity for the single-vortex test with $\mathrm{Re}=20,000$ and $\lambda=1$ on a uniform grid with $h=\frac{1}{1024}$. The region of each cell is filled by a single color that corresponds to the cell-averaged vorticity. No image smoothing is applied.
Following [37], we define an axisymmetric velocity field on $\Omega=[0,1]^{2}$ by
@@
where $r_{v}$ is the distance from the domain center $(\frac{1}{2},\frac{1}{2})^{T}$. The choices $R=0.2$ and $\Gamma=1$ give $\max(u_{\theta})=0.068$. A small viscosity $\nu=3.4\times 10^{-6}$ yields a high Reynolds number $\mathrm{Re}=20,000$. The initial condition of velocity is obtained by projecting cell averages of $\mathbf{u}$ in (70) ten times to make it approximately divergence-free.
Table 2: Errors and convergence rates of GePUP-ES-SDIRK for the single-vortex test with $\text{Re}=20,000$, $t_{0}=0.0$, $t_{e}=60$, $\lambda=1$ and the Courant number $\mathrm{Cr}=0.5$.
@@
(c) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$
@@
Figure 3: Evolutions of the modified energy, the SAV, and velocity divergence in the single-vortex test with $\mathrm{Re}=2\times 10^{4}$, $\mathrm{Cr}=0.5$, and $\lambda=1$. The abscissa in all subplots is time. In subplots (a) and (b), $h=\frac{1}{1024}$. In subplots (c) and (d), the curves marked by “square", “circle", and “diamond" represent the results for $h=\frac{1}{512}$, $\frac{1}{1024}$, and $\frac{1}{2048}$, respectively.
The tests are performed on four successively refined uniform grids. The time span $[0,60]$ is made long enough for the turbulent boundary layers to develop prominent Lagrangian coherent structures. Snapshots of the vorticity at time $t=40$ and at the final time $t=60$ are shown in Figure 2, where the essential features of vortex sheet roll-up and counter-vortices agree with those in [37].
As shown in Table 2, the convergence rates of the velocity in all norms are close to 4. In contrast, convergence rates of the scalar $q$ and its gradient have order reductions; this is due to the fact that the Neumann boundary condition in (54g) has to be obtained from spatial derivatives of the velocity and the calculation of these derivatives incurs order reductions in finite-volume discretizations.
Denote by $\mathbf{U}_{h}^{n}$ and $\mathbf{W}_{h}^{n}$ the finite-volume solutions that approximate cell averages of $\mathbf{u}$ and $\mathbf{w}$ at time $t^{n}$, respectively, and $r_{h}^{n}$ the computed value of the SAV $r$ at time $t^{n}$. The $L_{2}$-norm for a finite-volume solution $\mathbf{V}_{h}^{n}$ is defined as $\|\mathbf{V}^{n}_{h}\|_{\mathcal{C}}^{2}:=\sum_{\mathcal{C}_{\mathbf{i}}}\|\mathcal{C}_{\mathbf{i}}\|\cdot|\mathbf{V}^{n}_{h,\mathbf{i}}|^{2}$ with $\mathcal{C}_{\mathbf{i}}$ ranging over all control volumes. The modified energy is then computed as $\mathcal{E}_{h}^{n}:=\frac{1}{2}\left(\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}+|r_{h}^{n}|^{2}\right)$.
As shown in Figures 3(a,b), $\mathcal{E}_{h}^{n}$ decreases monotonically over the entire simulation with $|r_{h}^{n}-1|$ remaining around $10^{-11}$, indicating that the (unmodified) kinetic energy $\frac{1}{2}\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}$ also decreases monotonically.
In contrast, the $L_{2}$-norm of velocity divergence on the coarsest grid $h=\frac{1}{512}$ shown in Figures 3(c,d) first decreases dramatically during the first several seconds, then gradually increases, and oscillates within a certain range. Furthermore, the range of oscillation decreases quickly as the grid is refined. To understand this evolution pattern, we consider the fully discrete counterpart of (58), i.e.,
@@
where $\mathbf{D}$ represents the discrete divergence, $\mathbf{G}$ the discrete gradient, and $O(h^{p})$ errors of spatial discretization. Our numerical results, such as those in Figure 3, suggest $p>1$.

### v2

@@ 5.2 Semi-implicit GePUP-ES-SDIRK schemes
In light of Theorems 5.2 and 5.4, one way to preserve the monotonic decrease of the kinetic energy and the exponential decay of the divergence residue is to employ an algebraically stable RK method in GePUP-ES-RK. Gauss-Legendre RK methods are algebraically stable and have a minimal number of stages for a given temporal order of accuracy. However, their employment in GePUP-ES-RK necessitates either the coupling of all intermediate stage values of $\mathbf{w}$ or the use of complex arithmetic. Thus we turn to singly diagonal implicit RK (SDIRK) methods that satisfy
@@
Definition 5.5
A GePUP-ES-SDIRK (GES) scheme is a GePUP-ES-RK scheme where an algebraically stable SDIRK method is employed as the RK method and stage values for each intermediate stage $i=1,\ldots,s$ are decomposed as
@@
    where $\widetilde{\mathbf{w}}^{(i)}$ is an approximation of $\mathbf{w}^{(i)}$ obtained by (50) and (EAV-1,2).
  2. (GES.2)
@@
    $\left\{\begin{array}[]{rl}(1-\nu\gamma k\Delta)\mathbf{w}_{0}^{(i)}\par&=\mathbf{u}^{n}+k\sum\limits_{j=1}^{i}a_{i,j}\left(\mathbf{g}^{(j)}-\nabla q_{1}^{(j)}\right)+\nu k\sum\limits_{j=1}^{i-1}a_{i,j}\Delta\mathbf{w}_{0}^{(j)};\\ (1-\nu\gamma k\Delta)\mathbf{w}_{\ell}^{(i)}&=-ka_{i,\ell}\left(\widetilde{\mathbf{u}}^{(\ell)}\cdot\nabla\widetilde{\mathbf{u}}^{(\ell)}+\nabla q_{2}^{(\ell)}\right)+\nu k\sum\limits_{j=\ell}^{i-1}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)}.\end{array}\right$.
  4. (GES.4)
@@
The above steps (GES.1–6) are direct consequences of the GePUP-ES-RK scheme (48), the property (60) of SDIRK, and the decomposition (61).
Substitute the third decomposition in (61) into (49a,b), separate the terms with and without $r^{(i)}$, notice that $r^{(i)}$ is a scalar, and we have (GES.1–2). In step (GES.1), we decouple $q^{(i)}$ from $\mathbf{w}^{(i)}$ by approximating $\mathbf{w}^{(i)}$ with $\widetilde{\mathbf{w}}^{(i)}$ using the same method for calculating $\widetilde{\mathbf{u}}^{(i)}$.
For SDIRK, the upper bound of the summation in (49d) is $i$. Substitute (61) into (49c,d), separate the terms with and without $r^{(i)}$, switch the summation order by $\sum_{j=1}^{i}a_{i,j}\sum_{\ell=1}^{j}r^{(\ell)}\Delta\mathbf{w}_{\ell}^{(j)}=\sum_{\ell=1}^{i}r^{(\ell)}\sum_{j=\ell}^{i}a_{i,j}\Delta\mathbf{w}_{\ell}^{(j)}$, and we have (GES.3). Then (GES.4) follows from (49g).
(GES.5) follows from substituting the second equation in (61) into (49f),
@@
and moving the $r^{(i)}$-term on the RHS to the LHS. Finally, the parenthesis on the LHS of (GES.5) must be no less than 1 because of (46), Lemma 3.10, and the inner product of $\mathbf{u}_{i}^{(i)}$ to the equation on $\mathbf{w}_{i}^{(i)}$ in (GES.3).
(48) and (GES.1–6) are the complete algorithmic steps of the GES scheme. Theorems 5.2 and 5.4 state that both the velocity divergence and the modified energy in GES decrease monotonically provided that the employed SDIRK is algebraically stable; an example is given in (62).
In this work, we discretize the continuous spatial operators in Definition 5.5 by the fourth-order collocated finite-volume operators in (zhang16:_GePUP, , Sec. 3 & 4) to obtain a fully discrete GES scheme. However, it is emphasized that what we have proposed is not a single scheme but a space of solvers, each of which can be easily constructed by making menu choices for ‘orthogonal’ policies that span the solver space. The orthogonal structure of this solver space is very conducive to reusing in a black-box manner the legacy of classical finite volume/difference methods and the wealth of theory and algorithms for numerically solving ODEs; see Table 1. It is the GePUP-E formulation that makes this black-box reuse possible.
It would be ideal if the conclusions of Theorems 5.2 and 5.4 could also hold in the fully discrete case. Unfortunately, for the finite volume discretization on collocated grids, the fully discrete counterparts of Theorems 5.2 and 5.4 only hold asymptotically, i.e., in the limit of $k$ and $h$ simultaneously approaching zero. We defer to future research the investigation of suitable spatial discretizations so that Theorems 5.2 and 5.4 also hold in the fully discrete case.
@@ 6 Tests
In this section, we test a fully discrete GES scheme by several numerical experiments to confirm the analytic results in previous sections. We employ the fourth-order finite-volume discretizations in zhang16:_GePUP and a fourth-order, algebraically stable SDIRK method proposed by Du, Ju & Lu du19:sdirk ,
$\begin{tabular}[h]{c|ccc}$\gamma$&$\gamma$&&\\ $\frac{1}{2}$&$\frac{1}{2}-\gamma$&$\gamma$&\\ $1-\gamma$&$2\gamma$&$1-4\gamma$&$\gamma$\\ \hline\cr&$\mu$&$1-2\mu$&$\mu$\end{tabular},\qquad\left\{\begin{array}[]{l}\gamma=\frac{1}{\sqrt{3}}\cos\frac{\pi}{18}+\frac{1}{2},\\ \mu=\frac{1}{6(2\gamma-1)^{2}}.\end{array}\right$. (62)
Table 1: Choices of the main orthogonal policies that constitute the tested GES scheme as a fully discrete GePUP-ES-RK.
Orthogonal policies For the tested GES scheme Other possible options
@@
variable location collocated (as in zhang16:_GePUP ) staggered
@@
in Definitions 5.1 and 5.5 (50) & (EAV-1,2) both in time and in space
Along with other possible options of the major orthogonal policies, we show in Table 1 our choices that constitute the particular GES scheme to be tested. According to the analysis in Section 5, this particular INSE solver should be fourth-order accurate both in time and in space, preserve monotonic decrease of the total kinetic energy, and have the velocity divergence well under control.
All numerical tests are performed on a rectangular domain in Cartesian coordinates, where the electric boundary conditions in (32b) are enforced by homogeneous Dirichlet conditions for the tangential velocity components and a homogeneous Neumann condition for the normal velocity. The initial cell-averaged velocity is calculated by Boole’s rule, a sixth-order formula of Newton-Cotes quadrature. Since exact solutions are unavailable, we define the computational error via Richardson extrapolation, i.e., by the difference of the solution on the current grid and that of the next finer grid.
For different values of the penalty parameter $\lambda=$1, 10, and 100, the corresponding error norms in all numerical tests remain the same up to the first two significant digits; these results differ from those reported in shirokoff11:_navier_stokes ; rosales21:_high_poiss_navier_stokes . However, this is not surprising since projecting $\mathbf{w}$ to $\mathbf{u}$ with $\mathbf{u}\cdot\mathbf{n}=0$ in (32c,d) and setting $\mathbf{w}^{0}=\mathbf{u}^{0}$ in (33) already imply a small magnitude of $\mathbf{n}\cdot\mathbf{w}$. Nonetheless, a positive-valued $\lambda$ guarantees the no-penetration condition be fulfilled. Hereafter we only show results in the case of $\lambda=1$.
@@ 6.1 Single-vortex tests
(a) tt=40
(b) tt=60
Figure 2: Snapshots of vorticity for the single-vortex test with $\mathrm{Re}=20,000$ and $\lambda=1$ on a uniform grid with $h=\frac{1}{1024}$ and $\mathrm{Cr}=0.5$. The region of each cell is filled by a single color that corresponds to the cell-averaged vorticity. No image smoothing is applied.
Following bch91:_second_order_projection_for_viscous_incompressible_flow , we define an axisymmetric velocity field on $\Omega=[0,1]^{2}$ by
@@
where $r_{v}$ is the distance from the domain center $(\frac{1}{2},\frac{1}{2})^{T}$. The choices $R=0.2$ and $\Gamma=1$ give $\max(u_{\theta})=0.068$. A small viscosity $\nu=3.4\times 10^{-6}$ yields a high Reynolds number $\mathrm{Re}=20,000$. The initial velocity is obtained by projecting cell averages of $\mathbf{u}$ in (63) ten times to make it approximately divergence-free.
Table 2: Errors and convergence rates of the GES scheme in Table 1 for solving the single-vortex test with Re = 20,000, $t_{0}=0.0$, $t_{e}=60$, $\lambda=1$ and the Courant number $\mathrm{Cr}=0.5$.
@@
(c) $\left\|\mathbf{D}\mathbf{W}_{h}^{n}\right\|_{\mathcal{C}}$ as in (64)
@@
Figure 3: Evolutions of the modified energy, the SAV, and the $L_{2}$-norm of $\mathbf{D}\mathbf{W}_{h}^{n}$ produced by the GES scheme in Table 1 for solving the single-vortex test with $\mathrm{Re}=2\times 10^{4}$, $\mathrm{Cr}=0.5$, and $\lambda=1$. The abscissa in all subplots is time. In (a) and (b), $h=\frac{1}{1024}$ and $k\approx 7.18\times 10^{-3}$. In (c) and (d), the curves marked by $\blacksquare$,” $\bullet$,” and $\blacklozenge$” represent the results for $h=\frac{1}{512}$, $\frac{1}{1024}$, and $\frac{1}{2048}$, respectively.
The tests are performed on four successively refined grids with uniform grid size $h$. The time span $[0,60]$ is made long enough for the turbulent boundary layers to develop prominent Lagrangian coherent structures. The uniform time step size $k$ is given by setting the Courant number $\mathrm{Cr}:=\frac{k}{h}\max(u_{\theta})$ at 0.5. Snapshots of the vorticity at time $t=40$ and at the final time $t=60$ are shown in Figure 2, where the essential features of vortex sheet roll-up and counter-vortices agree with those in bch91:_second_order_projection_for_viscous_incompressible_flow .
It is clear in Table 2 that convergence rates of the velocity are close to 4 in all norms while those of the scalar $q$ and its gradient show order reductions, which are caused by the fact that the Neumann boundary condition in (47g) has to be obtained from spatial derivatives of the velocity and calculating these derivatives incurs order reductions in finite-volume discretizations.
Let $\mathbf{U}_{h}^{n}$ and $\mathbf{W}_{h}^{n}$ denote finite-volume solutions that approximate cell averages of $\mathbf{u}$ and $\mathbf{w}$ at time $t^{n}$, respectively. Then the $L_{2}$-norm for a finite-volume solution $\mathbf{V}_{h}^{n}$ is defined as
$\begin{array}[]{l}\|\mathbf{V}^{n}_{h}\|_{\mathcal{C}}:=\sqrt{\sum_{\mathcal{C}_{\mathbf{i}}}\|\mathcal{C}_{\mathbf{i}}\|\cdot|\mathbf{V}^{n}_{h,\mathbf{i}}|^{2}}\end{array}$ (64)
where $\mathcal{C}_{\mathbf{i}}$ ranges over all control volumes. The modified energy is then
$\begin{array}[]{l}\mathcal{E}_{h}^{n}:=\frac{1}{2}\left(\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}+|r_{h}^{n}|^{2}\right),\end{array}$ (65)
where $r_{h}^{n}$ is the computed value of the SAV $r$ at time $t^{n}$.
As shown in Figure 3(a,b), over the entire simulation $\mathcal{E}_{h}^{n}$ decreases monotonically and $|r_{h}^{n}-1|$ remains below $2.0\times 10^{-11}$, indicating that the (unmodified) kinetic energy $\frac{1}{2}\|\mathbf{U}^{n}_{h}\|_{\mathcal{C}}^{2}$ also decreases monotonically.
In Figure 3(c,d), the $L_{2}$-norm of velocity divergence on the coarsest grid $h=\frac{1}{512}$ first decreases dramatically during the first several seconds, then gradually increases, and oscillates within a certain range. Furthermore, the range of oscillation decreases quickly as the grid is refined. To understand this evolution pattern, we consider the fully discrete counterpart of (51), i.e.,
@@
where $\mathbf{D}$ is the discrete divergence, $\mathbf{G}$ the discrete gradient, and $O(h^{p})$ errors of spatial discretization. Numerical results such as those in Table 2 show $p>1$.